# Training-free Camera Control for Video Generation

### v1

@@ 4.1 Experimental Settings
Our work is based on zeng2023PixelDance with 16 frames at resolution 256$\times$256, which is trained on public data. Note that our method could be incorporated with most video diffusion models to achieve camera control, thus we also include results performed on other backbones SVD ; luo2023videofusion . For fair comparison with other methods, we set all model’s DDIM steps as 25. Our method doesn’t require any additional training utilizing camera trajectories.
@@
We compare CamTrol with state-of-the-art works: AnimateDiff guo2023animatediff , MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , all three methods finetune on specific data. For quantitative evaluation, FVD unterthiner2018FVD , FID heusel2017FID and IS saito2020IS are used to assess video generation quality, while CLIPSIM wu2021CLIPSIM quantifies the similarity between generated video and input prompt. The evaluations are performed on 8 basic camera trajectories, details could be found in Appendix A.1. For evaluation dataset, we follow similar setting in MotionCtrl wang2023motionctrl , but with more complex text prompt. Concretely, we apply UCF-101 soomro2012UCF as the reference of calculating FID and FVD, where we randomly sample 125 prompts per camera motion type and form 1000 samples in total for assessment. We also conduct a user study to compare different methods with human evaluation. Details can be found in Appendix A.5.
@@ 4.2 Comparisons with State-of-the-art Methods
Quantitative evaluations are shown in Table  1. Compared to previous state-of-the-art work, CamTrol demonstrates superior performance across most quantitative metrics. As large-stride camera move sometimes causes deformation on objects, there appears little decrease in single frame’s quality, i.e. FID. While method like AnimateDiff guo2023animatediff exhibits better performance on this term, it takes as sacrifice the dynamics of generated video, which we will see in qualitative evaluations. For camera motion control, human evaluations from user study indicate that training-free CamTrol attains comparable performance even compared to other training-based methods, where the video diffusion models are heavily finetuned on specific camera motion data.
Table 1: Quantitative comparison. Experiments are conducted on 8 basic camera trajectories.
Method Metric User Study
[…]
AnimateDiff guo2023animatediff 3699.34 65.81 11.61 0.3073 1.12 0.82
MotionCtrl wang2023motionctrl 1768.25 78.26 14.60 0.3143 2.00 2.03
CameraCtrl he2024cameractrl 1383.68 76.18 14.04 0.3123 1.01 1.08
CamTrol 1134.71 69.70 15.87 0.3253 1.86 2.06
Figure 3: Qualitative comparison with finetune methods. CamTrol generates videos that adhere to desired camera movements, without compromising their dynamics and rationality.
@@
Qualitative comparison of different methods is illustrated in Fig. 3. All methods generate videos align well with desired camera motion, however, the constraints imposed by camera trajectories are sometimes too strong, that output videos have to compromise their quality and rationality to satisfy these movements. Some examples that demonstrate this are results of Zoom Out and Roll Anticlockwise, which looks more like flat images being zoomed or rolled on 2D plane instead of videos recorded in 3D spaces. The reason might lies on the limitations of proprietary dataset used for finetuning realestate10k ; yu2023mvimgnet , where most scenes are static, making pretrained model forget about knowledge of dynamic contents. In contrast, CamTrol makes full use of the powerful dynamics in base model and generate plausible videos with camera perspective change utilizing layout prior guidance, handles these circumstances gracefully. More comparisons at Appendix A.3.
@@ 4.3 Ablation Study
Figure 4: Comparisons with base model. Controlling camera motion via prompt engineering doesn’t work at most times. Instead, CamTrol offers robust control towards video’s camera movement in a training-free manner.
@@
To demonstrate that changes of camera motion are attributed to our method rather than the innate capability of video base model, we conduct ablation study to assess its effectiveness. We add prompts describing relevant camera moves such as camera rotates around it or camera zooms out and let video model understand by itself. The results are shown in Fig. 4. We can observe that even provided with prompts indicating how camera should move, base model fail to produce correct results. Instead, CamTrol  is able to achieve corredponding motion control without any instructions from input prompts.
@@
We employ ablation study to validate the effectiveness of layout prior guidance, and illustrate its necessity from two aspects: the completeness and dynamic of generated video. In Fig. 5, we showcase frames before and after noise prior guidance. With the position of camera changes, there appears regions unfilled in point cloud and causes blank spaces in rendered images (left part); Besides, due to the static nature of point cloud, rendered images remain stationary too (right part). Noise layout prior could compensate for these flaws, finally produce videos with inpainted vacancies and rationalized dynamics.
@@
Our proposed CamTrol can be seamlessly plugged and played with most pretrained video diffusion models to achieve training-free camera control. We test its performance on different base video models, including SVD SVD and VideoFusion luo2023videofusion , the results are illustrated in Fig. 8. Our approach remains effective applied to other video base models, demonstrating its strong robustness and generalization ability. More results with other base models can be found in Appendix A.4.
@@ 4.4 Further Applications
By combining different basic camera trajectories, CamTrol can support hybrid camera movements and endow generated video with cinematic charm. Besides hybrid control, benefit from explicit camera motion modeling in 3D space, CamTrol could support trajectories with precise coordinates, which means it can generate videos with any complicated motion. Results about hybrid and complex camera motions are shown in Fig. 1 and Appendix A.2.
@@
CamTrol supports camera movements at controllable scales. We provide some results in Fig. 7. By specifying different magnitudes of camera’s extrinsic matrix within point cloud spaces, rendered images will exhibit varying degrees of motion, leading to videos with distinct scales of camera movements. This further demonstrate the powerful controllability of CamTrol, and provides new pathway for video’s customized camera control.
@@ 5 Conclusion
In this paper, we propose a training-free and robust method CamTrol to offer camera control for off-the-shelf video diffusion models. It consists of two-stage procedure including explicit camera motion modeling in 3D point cloud space and video generation utilizing layout prior of noisy latents. Compared to previours work, CamTrol  does not require any additional finetuning on camera-annotated datasets, or self-supervised training via data augmentation, instead, it could be plugged and played with most video diffusion models to generate camera controllable videos with single image or text prompt as input. Comprehensive experimental results demonstrate the effectiveness and robustness of CamTrol for controlling camera motion for videos. Besides, we show that CamTrol produces impressive results in generating videos at different scales, with 3D rotations and more complicated trajectories.
@@ A.3 More Comparisons to State-of-the-Art Methods
A.4 More Results on other Base Video Models
@@
Our methods could be plugged and played in arbitrary video diffusion models to achieve camera control. We test its performance on two other video generation backbones including SVD SVD and VideoFusion luo2023videofusion , relevant results are shown in Fig. 16.
A.5 Details of User Study
In user study, we collect 7 text prompts for each of the 8 basic camera trajectories, yielding 56 samples in total. Users are asked to sort randomly displayed four videos generated by four methods separately, and rate them from the highest score to the lowest score (belongs to {3,2,1}, where higher score indicates better performance). The assess metrics including two aspects: alignment of video move with input camera motion, and video quality (such as the deformation, rationality and diversity) of generated contents during camera moves. After rating, we calculate the average score attained by each methods.
The prompts applied in user study are ensured with balance between different categories. We involve prompts from various contents and styles including natural scenes, human, animations, objects, animals and etc.
Prompts applied in user study are listed as below:
  1. 1.
    Blueberries were squeezed by fingers.
  2. 2.
    The bear dolls were dancing in a circle, with lights flashing behind them, Christmas atmosphere.
  3. 3.
    The wind is blowing, and the flames are gradually extinguishing.
  4. 4.
    A young hero sets out on a journey to find their missing parent, encountering friends and foes along the way.
  5. 5.
    A toy robot discovers it can move on its own and sets off on an adventure around the house.
  6. 6.
    A determined ant leads its colony on a journey to find a new home.
  7. 7.
    The Great Wall of China winds its way across rugged mountains and through lush valleys, a testament to ancient engineering.
  8. 8.
    The bear doll in the basket is jumping out from the basket.
  9. 9.
    My armor pieces kept fluctuating and turned into scale-like armor.
  10. 10.
    A little boy turned a page of the book.
  11. 11.
    Jellyfish swim in the deep sea.
  12. 12.
    The big tree growing by the sea, its leaves slowly turn yellow and begin to wither and fall.
  13. 13.
    At the top of a mountain, a hiker is greeted by a breathtaking view of snow-capped peaks and rolling valleys below.
  14. 14.
    A group of squirrels gather around a picnic blanket, nibbling on sandwiches and stealing sips of lemonade.
  15. 15.
    The road is changing strangely and disappearing and twisting.
  16. 16.
    Clouds and fog linger on the tops of mountains.
  17. 17.
    Easter eggs showed different expressions, some happy, some confused, some sad, and some angry.
  18. 18.
    A skilled dancer twirls gracefully across the stage, captivating the audience with every movement.
  19. 19.
    A curious cat discovers a hidden treasure in the backyard.
  20. 20.
    At the beach, children build sandcastles while seagulls swoop overhead, and families relax under colorful umbrellas.
  21. 21.
    In a bustling marketplace, vendors call out their wares as shoppers weave through the crowds, browsing colorful stalls.
  22. 22.
    On the busy streets, people come and go.
  23. 23.
    Sunlight filters through dense foliage in a tranquil grove.
  24. 24.
    A toy monkey sits at a pretend computer, typing away furiously as it pretends to work on important documents.
  25. 25.
    The Eiffel Tower stands tall against a backdrop of Parisian rooftops, its iron lattice illuminated by the setting sun.
  26. 26.
    The Colosseum in Rome stands as a majestic reminder of ancient gladiator battles and Roman history.
  27. 27.
    A cozy reading nook by a window.
  28. 28.
    A regal statue of a historical figure stands tall in a city square, gazing out over the bustling streets.
  29. 29.
    Let them perform a dance.
  30. 30.
    The goose is playing trumpet. slow motion. 2D animation.
  31. 31.
    The machine is making Spider-Man’s hood.
  32. 32.
    The girl danced in the water, splashing water and her hair fluttering the wind.
  33. 33.
    The man looked up at the peak.
  34. 34.
    A woman reading and taking notes.
  35. 35.
    A group of penguins are walking towards you.
  36. 36.
    A shimmering portal appears in the midst of a moonlit forest, leading to realms unknown.
  37. 37.
    In a bustling port, cranes swing cargo containers onto waiting ships, orchestrating a dance of logistics and trade.
  38. 38.
    At dawn, the sky ignites in a riot of colors over the tranquil expanse of a mist-covered lake.
  39. 39.
    A touch of enchanted ink brings the pages of an ancient spellbook to life, casting spells with each turn.
  40. 40.
    A travel blogger documents their journey around the world, sharing tales of exotic destinations and unforgettable experiences.
  41. 41.
    A playful puppy learns to fetch a stick for the first time.
  42. 42.
    In a cozy living room, friends gather around a crackling fireplace, sipping hot cocoa and sharing stories on a chilly evening.
  43. 43.
    A man fell asleep amid the flowers, roses swaying in the wind.
  44. 44.
    A tutorial video guides viewers through the step-by-step process of creating intricate origami designs.
  45. 45.
    Underneath a starry sky, the Northern Lights dance in shimmering curtains of green and purple, painting the night with magic.
  46. 46.
    A cluttered desk in a home office, with papers strewn about and a computer humming softly.
  47. 47.
    A towering skyscraper gleams in the sunlight, its sleek glass facade reflecting the surrounding cityscape.
  48. 48.
    A modernist home with clean lines and geometric shapes, surrounded by minimalist landscaping.
  49. 49.
    A children’s playground is a kaleidoscope of colors, with slides, swings, and climbing structures painted in playful shades.
  50. 50.
    Thunder and rain, ancient oriental dragons hover in the air.
  51. 51.
    A dog barking at you.
  52. 52.
    Multiple people raised their arms and struggled upward in the lava of hell.
  53. 53.
    An astronaut aboard a space station encounters a mysterious anomaly while conducting experiments.
  54. 54.
    A guitarist strums their instrument with gentle fingers, the melody flowing naturally from their movements.
  55. 55.
    A snow-capped mountain peak towers above a tranquil alpine lake, mirrored perfectly in its glassy surface.
  56. 56.
    A rainbow arcs across the sky after a refreshing rain shower, its vibrant colors painting a stunning backdrop against the clouds.

### v2

@@ 4.1 Experimental Settings
The major results presented in this paper are grounded on zeng2023PixelDance . To ensure a fair comparison with other state-of-the-art methods, we employ SVD as base model for quantitative evaluation. Inversion and generation steps among all methods are set to 25. Our method doesn’t require any additional training utilizing camera trajectories.
@@
We compare CamTrol with state-of-the-art works: AnimateDiff guo2023animatediff , MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , all three methods are finetune on SVD using specific data. In quantitative evaluation, FVD unterthiner2018FVD , FID heusel2017FID and IS saito2020IS are used to assess video generation quality, while CLIPSIM wu2021CLIPSIM quantifies the similarity between generated video and input prompt. With regard to the accuracy of camera motion, we adopt ParticleSFMzhao2022particlesfm and produce estimated camera trajectories from generated videos, with the use of Absolute Trajectory Error(ATE) measuring their differences compared to ground truth. Relative Pose Error(RPE) is calculated to assess between consecutive frames how well the relative motions match expected ones including their transition(RPE-T) and rotation part(RPE-R). Settings of evaluation dataset follow those established in MotionCtrl wang2023motionctrl , but include more complex trajectories. Specifically, we extract 51 distinct trajectories from RealEstate10krealestate10k , each paired with 10 prompts mentioned in MotionCtrl wang2023motionctrl , resulting in 510 samples in total for assessment. As AnimateDiff guo2023animatediff lacks the ability to handle complicated trajectories, it is only included in qualitative analysis part evaluating with 8 basic camera motions. Quantitative comparisons contain AnimateDiff guo2023animatediff can be found in Appendix A.3.
@@ 4.2 Comparisons with State-of-the-art Methods
Quantitative evaluations are shown in Table  1. Building upon the same base model as MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , our method demonstrates superior performance across all quantitative metrics concerning both video quality and camera motion accuracy. Despite being training-free, CamTrol outperforms those that rely heavily on extensive finetuning and large-scale annotated data, attaining the lowest score in ATE, RPE-T and RPE-R. Benefit from explicit camera movements modeling and motion inversion, our method is capable to produce videos align best with those tricky trajectories while maintaining their visual qualities.
Table 1: Quantitative comparisons.
Method Video Quality Motion Accuracy
FVD $\downarrow$ FID $\downarrow$ IS $\uparrow$ CLIP-SIM $\uparrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
MotionCtrl wang2023motionctrl 3576.40 239.10 7.58 0.2933 4.006 1.086 0.106
CameraCtrl he2024cameractrl 2922.99 243.98 8.07 0.2915 4.200 1.487 0.080
SVD+CamTrol 2832.59 227.36 8.07 0.3100 3.917 0.947 0.017
Figure 3: Qualitative comparisons with finetuned methods. CamTrol generates videos that adhere to desired camera movements, without compromising their dynamics and rationality.
@@
Qualitative comparisons of different methods are illustrated in Fig. 3. All methods generate videos align well with desired camera motion, however, the constraints imposed by camera trajectories are sometimes too strong, that output videos have to compromise their quality and rationality to satisfy these movements. Some examples that demonstrate this are results of Zoom Out and Roll Anticlockwise, which looks more like flat images being zoomed or rolled on 2D plane instead of videos recorded in 3D spaces. The reason might lies on the limitations of proprietary dataset used for finetuning realestate10k ; yu2023mvimgnet , where most scenes are static, making pretrained model forget about knowledge of dynamic contents. In contrast, CamTrol makes full use of the powerful dynamics in base model and generate plausible videos with camera perspective change utilizing layout prior guidance, handles these circumstances gracefully. More comparisons at Appendix A.3.
@@ 4.3 Ablation Study
Figure 4: Comparison with base model. Controlling camera motion via prompt engineering doesn’t work at most times. Instead, CamTrol offers robust control towards video’s camera movement in a training-free manner.
@@
To demonstrate that changes of camera motion are attributed to our method rather than the innate capability of video model, we conduct ablation study to assess its effectiveness. We add prompts describing certain camera moves(e.g. zooms out), letting video model understand by itself. The results are shown in Fig. 4. It could be observed that even provided with prompts indicating how camera should move, base model fail to produce correct results. Instead, CamTrol  is able to implement designated motion control without any instructions from text prompts.
@@
We employ ablation study to validate the effectiveness of layout prior guidance, illustrating its necessity from two aspects: the completeness of vacancies and dynamic of generated video. In Fig. 5, we showcase frames before and after noise prior guidance. With camera pose changes, there appears regions unfilled in point cloud and causes blank spaces in rendered images(left part); Besides, due to the static nature of point cloud, rendered images remain stationary(right part). Noise layout prior could compensate for these flaws, finally produce videos with inpainted vacancies and rationalized dynamics.
@@
Our proposed CamTrol can be seamlessly plugged and played with most pretrained video diffusion models to achieve training-free camera control. We present visual results showcasing its applications based on different video base models, including SVD SVD and VideoFusion luo2023videofusion , in Fig. 8. Our approach remains effective applied to alternative video base models, demonstrating its strong robustness and generalization ability. More results with other base models can be found in Appendix A.5.
@@ 4.4 Further Applications
By combining different basic camera trajectories, CamTrol can support hybrid camera movements and endow generated video with cinematic charm. Besides this, explicit motion modeling also equips CamTrol with the abilities to support trajectories containing precise coordinates, which means it can generate videos presenting any complicated camera movement. Results about hybrid and complex motions are shown in Fig. 1 and Appendix A.2.
@@
CamTrol supports camera movements at controllable scales. We provide some results in Fig. 7. By specifying different magnitudes of camera’s extrinsic matrix within point cloud spaces, rendered images will exhibit varying degrees of motion, leading to videos with distinct scales of camera movements. This further demonstrate the powerful controllability of CamTrol, and provides a new pathway for video’s customized camera control.
@@ 5 Conclusion
In this paper, we propose a training-free and robust method CamTrol to offer camera control for off-the-shelf video diffusion models. It consists of two-stage procedure including explicit camera motion modeling in 3D point cloud space and video generation utilizing layout prior of noisy latents. Compared to previours work, CamTrol  does not require any additional finetuning on camera-annotated datasets, or self-supervised training via data augmentation, instead, it could be plugged and played with most video diffusion models to generate camera controllable videos with single image or text prompt as input. Comprehensive experimental results demonstrate the effectiveness and robustness of CamTrol for controlling camera motion for videos. Besides, we show that CamTrol produces impressive results in generating videos with 3D rotations, complicated trajectories, and different moving scales.
@@ A.3 More Comparisons to State-of-the-Art Methods
We also present quantitative comparisons that include AnimateDiff guo2023animatediff as part of the evaluation. As AnimateDiff guo2023animatediff lacks the ability to handle complicated trajectories, experiments are conducted focusing on 8 basic trajectories described in A.1. To expand the diversity of the models evaluated, we employ CamTrol on zeng2023PixelDance for assessment. With regard to basic trajectories, we apply UCF-101 soomro2012UCF as the reference of calculating FID and FVD, from where we randomly sample 125 prompts per camera motion type and form 1000 samples in total for assessment. Relevant results are exhibited in Table 3.
Table 3: Quantitative comparisons on basic trajectories.
Method Video Quality
FVD $\downarrow$ FID $\downarrow$ IS $\uparrow$ CLIP-SIM $\uparrow$
AnimateDiff guo2023animatediff 1582.93 73.79 13.39 0.3196
MotionCtrl wang2023motionctrl 1768.25 78.26 14.60 0.3143
CameraCtrl he2024cameractrl 1383.68 76.18 14.04 0.3123
CamTrol+zeng2023PixelDance 1134.71 69.70 15.87 0.3253
A.4 Ablation on Depth Optimization
Ablation results relevant to the effectiveness of depth optimization are presented in Table 4. While it affects little on CLIPSIM, depth optimization exhibits benefits on other video quality assessments(FVD, FID, IS) and camera trajectory criteria(ATE, RPE). Experiments are conducted on complicated 10 trajectories extracted from RealEstate10k realestate10k , each with 10 prompts mentioned in MotionCtrl wang2023motionctrl . We assume that depth optimization offers better alignment between adjacent views rendered from point cloud, which benefits both perceptual quality and camera motion accuracy of generated videos.
Table 4: Ablation on depth coefficient optimization.
[…]
w/o optimization 5033.44 219.01 5.67 0.2949 4.093 1.131 0.046
w/ optimization 4977.88 218.36 5.72 0.2947 3.906 0.978 0.049
A.5 More Results on other Base Video Models
@@
Our methods could be plugged and played with arbitrary video diffusion models to achieve camera control. We showcase its performance based on two alternative video generation backbones including SVD SVD and VideoFusion luo2023videofusion , relevant results are exhibited in Fig. 16.